OpenStudy (anonymous):

Fraction question I have 69/h and I was to isolate h, will it matter if i multiply or divide by 69 to isolate h?

zepdrix (zepdrix):

Your problem is 69/h...? There is nothing else in the problem? Because I don't see how you're suppose to isolate h if we can't manipulate the other side of the equation D: hmm

OpenStudy (anonymous):

its a word problem and i have the half life equation and i'm trying to find the half life which is h, so far i have log1/square root 2/log 1/2 =69/h

zepdrix (zepdrix):

I can't understand with the way you wrote that... Is this what it looks like?\[\huge \frac{\log\left(\frac{1}{\sqrt2}\right)}{\log\left(\frac{1}{2}\right)}=\frac{69}{h}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Ummm the way I would do this, might be a little confusing. It might be easier if you remember how to do Cross Multiplication.

zepdrix (zepdrix):

Anyway this is how I would do it :D maybe it'll make sense. Start by multiplying both sides by h.\[\huge h\cdot \frac{\log\left(\frac{1}{\sqrt2}\right)}{\log\left(\frac{1}{2}\right)}=\frac{69}{h}\cdot h\]The h's on the right will cancel out, giving us,\[\huge h\cdot \frac{\log\left(\frac{1}{\sqrt2}\right)}{\log\left(\frac{1}{2}\right)}=69\]Then to get rid of that big fraction, multiply both sides by it's reciprocal (the same fraction but FLIPPED).\[\huge h=69\frac{\log\left(\frac{1}{2}\right)}{\log\left(\frac{1}{\sqrt2}\right)}\] It's just a matter of moving things around to isolate h. If that was rather confusing, think back to cross multiplication :D

OpenStudy (anonymous):

what if h was on top, so h/69, will it still be the same result, because thats the way i'm thinking of it, but im multiplying both sides by the reciprocal or fraction to isolate h

zepdrix (zepdrix):

if h was on top, we would have a lot less work, we only have to move the 69 to the other side. So we would have ended up multiplying both sides by 69, and that's all.

OpenStudy (anonymous):

ok, thanks